# A new indeterminacy-based quantum theory

### v4

@@ Acknowledgments
So, we have here familiar situations in the philosophy of mind. Alice could, in fact, have experiences in these situations, and EnDQT can account for them. There is much more to say about this. Future work will go into more detail on a), b), and c). Note that a), b), and c) are options that may be adopted if we reject the absoluteness of observed events assumption in the way EnDQT did (section 3).
@@
It’s important to notice that if we consider realist Wigner’s friend scenarios, the position adopted by EnDQT regarding the friend’s experiences and the adoption of the above hypotheses shouldn’t be seen as something restricted to EnDQT. More concretely, if extended Wigner’s friend scenarios become realizable one day, it will very likely be via quantum computers and quantum agents running on those quantum computers as friends instead of human friends (see Wiseman et al., 2023 for a proposal). Assuming the controversial position that such quantum agents have mental content, which is a requirement if we want this version to mimic the original extended Wigner’s friend version, many realist interpretations of QT will be pressed to assume that quantum agents don’t have internally determinate mental content. This is because, plausibly, [^80] I am setting aside strong emergentist and dualist perspectives about such content here. their experiences will depend on superpositions of qubits. As it is recognized by many MWI proponents, [^81] See most prominently, Wallace (2012, section 10.3). we can have robust branching into worlds when there is decoherence, but inside some quantum computers, we shouldn’t often have such branching because there isn’t a lot of decoherence (at least ideally and in many architectures of quantum computers). Many proponents of interpretations such as the MWI won’t consider that, in many situations, there is enough robust branching inside the quantum computer so that we could have something like an agent with determinate mental content running on those circuits. Spontaneous collapse theories won’t also consider that there is such an agent because they don’t consider that collapses happen (at least frequently) in situations like those within a quantum computer.
@@
Wiseman et al. (2023) basically acknowledge the above in the case of spontaneous-collapse theories, saying that the "thoughts [of the artificial agent] are thus not real in the way that my thoughts as a human are real." This amounts to the rejection of the "Friendliness" assumption of the theorem of Wiseman et al. (2023). My claim is that EnDQT also rejects this assumption, as well as the (at least some influential versions of the) MWI. [^82] Of course, accepting b) and c), one shouldn’t talk in terms of the reality of the thoughts/mental content. Instead, we should talk in terms of how different they are from our thoughts because the quantum agents have thoughts, they are just different from what we typically conceive our thoughts to be.
@@
Furthermore, although single-world relationalists can account for the relative friend’s experiences and prima facie this is an advantage relative to EnDQT, there is a good case to be made that these experiences aren’t absolute. A more careful inspection of single-world relationalist views, such as Relational Quantum Mechanics, [^83] See, e.g., Rovelli (1996) and Di Biagio & Rovelli (2021). shows that relative to some systems, other systems’ mental content can be indeterminate since relative to one system, the other system might be in a superposition of quantum states that the mental content depends on. So, views such as Relational Quantum Mechanics, in these circumstances, would be in a similar position as EnDQT and be subject to a version of the above worry.
Appendix B: The basics of an ontology of quantum properties
@@
and the observable $S_{z}$ that acts on the Hilbert space of $S$, represents the quantum property spin-z of $S$. The spin-z of $S$ has a degree of differentiation $D^{*}=0$ and we consider that the system has an undifferentiated spin-z. This is because this system is not interacting with any other systems (note that it would still have an undifferentiatied spin if it interacted with systems that don’t have the DC).
Let’s consider a system $E$, constituted by many subsystems that are interacting with $S$. For instance, $S$ with quantum properties spin in different directions that interacts strongly with the many systems also with spin in multiple directions that constitute $E$. I will again put a subscript SDC in the systems that belong to an SDC. A system belonging to an SDC will have a stably differentiated quantum property represented via its quantum states when it interacts, in agreement with the CDCs (Section 2), with another system $S$, decohering it, and thus giving rise to interactions belonging to an SDC. So, if $S$ is interacting with a system $E$ belonging to an SDC and having the DC-S (Section 2), we have that
@@
a measure of the degree of differentiation of the different $D^{\*}$-P of $S$ in spacetime region over time $t$ for the simple scenarios that we are considering will be given by the von Neumann entropy [^84] Given a density operator $\rho_{S}$ for quantum system $S$, the von Neumann entropy is $S\left(\rho_{S}\right)=-\operatorname{tr}\left(\rho_{S}\ln\rho_{S}\right)$. $S\left(\hat{\rho}_{S}\right)$ is zero for pure states and equal to $\ln N$ for maximally mixed states in this finite-dimensional case. $S\left(\hat{\rho}_{S}(t)\right)$ of $\hat{\rho}_{S}(t)$ over $\ln N$, where $N$ is the number of eigenvalues of $\hat{\rho}_{S}(t)$,
@@
Thus, we can measure and represent the degree of differentiation $\mathrm{D}^{*}$ of the quantum property D*-P that $S$ will end up with at the end of the interaction with $E$ at $\mathrm{t}$ with $0\leq$ $D^{*}(P,S,t)\leq 1$, and the differentiation timescale (which is inferred via the decoherence timescale).
@@
I will come back below to the intuition for why we can consider that there are values with degrees of determinacy. For now, note that not all interactions with a system give rise to systems having a determinate value, although there is something that changes in the quantum properties of the systems under these interactions. As a toy example, consider the spin of a particle in different directions in a series of Stern-Gerlach devices without letting the particles hit a screen between each device. The inhomogeneous magnetic field leads the subsystem of the particle with a spin in a certain direction to interact with the subsystem with the quantum property position, leading to their entanglement. There is something that changes in the spin direction of the quantum systems when the particle goes from one magnet to the other, but there is no determinate value arising. If there was, we would have an irreversible process, and thus, we wouldn’t be able to reverse the result of the operations of the magnet via a Stern-Gerlach interferometer. So, the spin of the system that interacts with the other subsystem of the particle has an indeterminate value, although there is something that changes in the quantum property that corresponds to this indeterminate value.
@@
As I have mentioned in Section 2, pragmatic reversible decoherence models allow us to infer and represent $E$ and $S$ interacting but having indeterminate values. This occurs when $E$ doesn’t belong to an SDC, not having the DC. The Stern-Gerlach case above is a case appropriately modeled via a pragmatic irreversible decoherence model. [^85] See, e.g., de Oliveira & Caldeira (2006) for one such model. I will call the interactions represented and inferred via the pragmatic reversible decoherence models, unstable differentiation interactions. During these interactions, both systems continue with their quantum properties undifferentiated.
@@
As I have been assuming above, it is plausible to consider that some quantum properties can be stably differentiated to a certain degree, and this impacts the subsequent degree of determinacy of the value that arises from a quantum property. Let’s look at the intuition for this. In the double-slit experiment, if the detectors at the slits interact with a quantum system weakly in such a way that we can’t fully distinguish in which slit it passed we get some disappearance of interference. These interactions will give rise to a low entanglement between the position and the degrees of freedom of the detector. Furthermore, the more these interactions distinguish the path of the system, the more entanglement we have between the position of the target system and the degrees of freedom of the detector, and the more the interference disappears until it disappears completely under maximal entanglement. So, I have considered that stable differentiation of a quantum property comes in degrees, as well as the determinacy of the resultant values.
To explain the dependence between the degree of determinacy of values of systems and the degree of differentiation of their quantum properties, I will adopt a functionalist account of indeterminacy. Very roughly, functionalism about property $\mathrm{P^{*}}$ is the position that $\mathrm{P^{*}}$ is the property of having some other property P in a certain situation or having specific features. The functionalist position provides an account of the dependence relation between the values properties (henceforward, values) $\mathrm{v}$ that I have been talking about, which come in terms of different degrees of determinacy and quantum properties. We have that
@@
Pipa (2024) will enter into further details about this ontology. It may initially seem pedantic compared with the simpler ontology of flashes and observables. However, it captures more structure represented by quantum states (and decoherence) than the flashes. Systems don’t only have determinate values under interactions (which would be analogous to the flashes), they have quantum properties with different degrees of differentiation that change over time.
Appendix C: Future directions
Future work should seek to test and further develop EnDQT. As one can see, EnDQT has a series of distinct features when compared with other quantum theories. At first, it seems that it will be very hard to distinguish EnDQT empirically from the other unitary interpretations of QT because, in practice like EnDQT, all of them appeal to (irreversible) decoherence connected with some environments in one way or another.
However, I have mentioned in section 2 several distinct predictions of EnDQT. Furthermore, since EnDQT doesn’t violate relativistic causality, we can regard that as indirect evidence for this view since such violations have yet to be seen. Also, EnDQT offers a finer account of how determinacy propagates than other views since certain interactions between systems become important. If this finer account is further developed and empirically confirmed, it provides good support for EnDQT since the other interpretations of QT don’t require it. So, first, we should find ways to test the CDCs with their distinct predictions discussed in section 2 and possibly propose and test new ones.
Another way to find confirmatory evidence for EnDQT is by searching for other phenomena that it can further explain. Future work should seek to investigate SDCs associated with spacetime and gravity in order to see if EnDQT could help achieve the integration of QT with gravity. Indeed, SDCs and gravity/spacetime have some aspects in common. It was shown via the Hawking-King-McCarthy-Malament theorem (Hawking et al., 1976; Malament, 1977) that (roughly) the causal order of relativistic spacetimes $determines$ its geometry up to a conformal factor. A $causal$ order of events also arises between the non-spacelike separated events that constitute SDCs. If we consider the SDC, $X\rightarrow$ $Y\rightarrow Z$, we obtain the following "causal" order of events, $E_{X\rightarrow Y}\rightarrow E_{Y\rightarrow Z}.E_{X\rightarrow Y}$ are the events occurring when $X$ gives rise to $Y$ having a determinate value and transmits the $\mathrm{DC}$ to $Y$, and $E_{Y\rightarrow Z}$ are the events occurring when $Y$ gives rise to $Z$ having a determinate value and transmits the DC to $Z$. More generally, we can have a DAG representing the events where a system transmits the DC to another one or potentially does so, and it also gives rise to each other having determinate values. Given the SDCs-starting hypothesis, the beginning of SDCs like the origin of gravity and spacetime may date back to the beginning of the universe, both SDCs and spacetime expand and they are widespread. Perhaps, the events that arise from EnDQT could even give rise to its own version of the causal set-like structures [^86] Bombelli et al. (1987). In this case, the first events that constitute a set (i.e., the so-called ”post”) could arise from initiators. These events might allow us to hypothesize initiators that don’t involve the inflaton. Such events could be the ones behind the origin of spacetime and gravity. because SDCs seem to naturally give to the DAGs that characterize such structures, but where such sets naturally arise from the quantum level contrary to the causal sets program.
@@
In this appendix, I put into practice some of the above features of EnDQT to see how it can account for interference phenomena via a simple example. [^87] This model is based on von der Linde (2021). See Cohen-Tannoudji et al. (1997) for a more extensive discussion of this framework.
@@
represent the number of photons (the particle number) in the channel $\mathrm{Ch}$, where each channel is associated with a radiation mode.
@@
After BS2, the system will interact with the detector D2, giving rise to a system having a 1 particle number determinate value during the interaction. Note that it’s assumed that D2 is connected to an SDC.
Let’s now clarify how interactions that lead to a determinate value work by examining what happens when detector D3 is placed at B1 (see Figure 5). This detector interacts with the quantum system, annihilating the above interference phenomenon. I am going to adopt the same SDC subscripts convention that I have adopted in the last sections. The interactions at time t’ involving D3 (and omitting the interactions with D1 and D2) lead to the following state,
@@
where these interactions can be represented via decoherence models. [^88] For simplicity, I will not analyze in detail decoherence in the Fock basis and assume that the Schrödinger picture is applicable. See Walls and Milburn (1995), Mcclung et al. (2010), and Myatt et al. (2000) for a detailed account. I will also assume that a notion of spatiotemporal localization of particles arises during these interactions. See Fraser (2022) for a survey of different options that consider particles as non-fundamental, but emergent. How does EnDQT interpret the above phenomenon? First, note that contrary to $\left|E_{1}(t)\right\rangle,\left|E_{0}(t)\right\rangle$ concerns the inexistence of the measurement signal. It will also mean that the measurement device interacted with a subsystem, giving rise locally to a 0 particle number determinate value (i.e., the vacuum).
@@
Let’s now consider instead the situation where the detectors are isolated from interacting with elements of an SDC, not belonging to an SDC as well. The quantum state $\mid$ Final $\left.{}^{\prime}\left(t^{\prime}\right)\right\rangle$ isn’t anymore applicable to correctly represent the situation inside the lab. We would rather have

### v5

@@ Acknowledgments
So, we have here familiar situations in the philosophy of mind, which involves externalism. Alice could, in fact, have experiences in these situations, and EnDQT can account for them. There is much more to say about this. Future work will go into more detail on a), b), and c). Note that a), b), and c) are options that may be adopted if we reject the absoluteness of observed events assumption in the way EnDQT did (Section 3).
@@
It is important to notice that if we consider realist Wigner’s friend scenarios, the position adopted by EnDQT regarding the friend’s experiences and the adoption of the above hypotheses shouldn’t be seen as something restricted to EnDQT. More concretely, if extended Wigner’s friend scenarios become realizable one day, it will very likely be via quantum computers and quantum agents running on those quantum computers as friends instead of human friends (see [103] for a proposal). Assuming the controversial position that such quantum agents have mental content, which is a requirement if we want this version to mimic the original extended Wigner’s friend version, many realist interpretations of QT will be pressed to assume that quantum agents don’t have internally determinate mental content. This is because, plausibly,133133133I am setting aside strong emergentist and dualist perspectives about such content here. their experiences will depend on superpositions of qubits. As it is recognized by many MWI proponents,134134134See most prominently, ([95], Section 10.3). we can have robust branching into worlds when there is decoherence, but inside some quantum computers, we shouldn’t often have such branching because there isn’t a lot of decoherence (at least ideally). Many proponents of interpretations such as the MWI won’t consider that, in many situations, there is enough robust branching inside the quantum computer so that we could have something like an agent with determinate mental content running on those circuits. Spontaneous collapse theories won’t also consider that there is such an agent because they don’t consider that collapses happen (at least frequently) in situations like these ones within a functional quantum computer.
@@
[103] basically acknowledge the above in the case of spontaneous-collapse theories, saying that the “thoughts [of the artificial agent] are thus not real in the way that my thoughts as a human are real.” This amounts to the rejection of the “Friendliness” assumption of the theorem of [103]. My claim is that EnDQT also rejects this assumption, as well as the (at least some influential versions of the) MWI.135135135Of course, accepting b) and c), one shouldn’t talk in terms of the reality of the thoughts/mental content. Instead, we should talk in terms of how different they are from our thoughts because the quantum agents have thoughts, they are just different from what we typically conceive our thoughts to be.
@@
Furthermore, although single-world relationalists can account for the relative friend’s experiences and prima facie this is an advantage relative to EnDQT, there is a good case to be made that these experiences aren’t absolute. A more careful inspection of single-world relationalist views, such as Relational Quantum Mechanics,136136136See, e.g., [88, 32]. shows that relative to some systems, other systems’ mental content can be indeterminate since relative to one system, the other system might be in a superposition of quantum states that the mental content depends on. So, views such as Relational Quantum Mechanics, in these circumstances, would be in a similar position as EnDQT and be subject to a version of the above worry.
Appendix C: The basics of an ontology of quantum properties
@@
and the observable $S_{z}$ that acts on the Hilbert space of $S$, represents the quantum property spin-z of $S$. The spin-z of $S$ has a degree of differentiation $D^{*}=0$ and we consider that the system has an undifferentiated spin-z. This is because this system is not interacting with any other systems (note that it would still have an undifferentiated spin if it interacted with systems that don’t have the DC).
Let’s consider a system $E$, constituted by many subsystems that are interacting with $S$. For instance, $S$ with quantum properties spin in different directions that interacts strongly with the many systems, also with spin in multiple directions, that constitute $E$. I will again put a subscript SDC in the systems that belong to an SDC. A system belonging to an SDC will have a stably differentiated quantum property represented via its quantum states when it interacts, in agreement with the CDCs (Section 2), with another system $S$, decohering it and thus giving rise to interactions belonging to an SDC. So, if $S$ is interacting with a system $E$ belonging to an SDC and having the DC-$S$ (Section 2), we have that
@@
a measure of the degree of differentiation of the different $D^{\*}$-P of $S$ in spacetime region over time $t$ for the simple scenarios that we are considering will be given by the von Neumann entropy137137137Given a density operator $\rho_{S}$ for quantum system $S$, the von Neumann entropy is $S\left(\rho_{S}\right)=-\operatorname{tr}\left(\rho_{S}\ln\rho_{S}\right)$. $S\left(\hat{\rho}_{S}\right)$ is zero for pure states and equal to $\ln N$ for maximally mixed states in this finite-dimensional case. $S\left(\hat{\rho}_{S}(t)\right)$ of $\hat{\rho}_{S}(t)$ over $\ln N$, where $N$ is the number of eigenvalues of $\hat{\rho}_{S}(t)$,
@@
Thus, we can measure and represent the degree of differentiation $\mathrm{D}^{*}$ of the quantum property D*-P that $S$ will have at the end of the interaction with $E$ at $\mathrm{t}$ with $0\leq$ $D^{*}(P,S,t)\leq 1$, and the differentiation timescale (inferred from the decoherence timescale).
@@
I will return below to the intuition for why we can consider that there are values with degrees of determinacy. For now, note that not all interactions with a system give rise to systems having a determinate value, although there is something that changes in the quantum properties of the systems under these interactions. As a toy example, consider the spin of a particle in different directions in a series of Stern-Gerlach devices without letting the particles hit a screen between each device. The inhomogeneous magnetic field leads the subsystem of the particle with a spin in a certain direction to interact with the subsystem with the quantum property position, leading to their entanglement. Something changes in the spin direction of the quantum systems when the particle goes from one magnet to another, but no determinate value arises. If there was, we would have an irreversible process, and thus, we wouldn’t be able to reverse the result of the operations of the magnet via a Stern-Gerlach interferometer. So, the spin of the system that interacts with the other subsystem of the particle has an indeterminate value, although there is something that changes in the quantum property that corresponds to this indeterminate value.
@@
As I have mentioned in Section 2, pragmatic reversible decoherence models allow us to infer and represent $E$ and $S$ interacting but having indeterminate values. This occurs when $E$ doesn’t belong to an SDC, not having the DC. The Stern-Gerlach case above is a case appropriately modeled via a pragmatic irreversible decoherence model.138138138See, e.g., [31] for one such model. I will call the interactions represented and inferred via the pragmatic reversible decoherence models, unstable differentiation interactions. During these interactions, both systems continue with their quantum properties undifferentiated.
@@
As I have been assuming above, it is plausible to consider that some quantum properties can be stably differentiated to a certain degree, and this impacts the subsequent degree of determinacy of the value that arises from a quantum property. Let’s look at the intuition for this. In the double-slit experiment, if the detectors at the slits interact with a quantum system weakly in such a way that we cannot fully distinguish in which slit it passed we get some disappearance of interference. These interactions will give rise to a low entanglement between the position and the degrees of freedom of the detector. Furthermore, the more these interactions distinguish the path of the system, the more entanglement we have between the position of the target system and the degrees of freedom of the detector, and the more the interference disappears until it disappears completely under maximal entanglement. So, I have considered that stable differentiation of a quantum property comes in degrees and the determinacy of the resultant values.
Differentiation and determinacy are related and this will allow us to provide an analysis of quantum indeterminacy and determinacy. This relation will establish that property $P^{*}$, in this case, a value property, is the property of having some other property P having specific features. I will thus consider that
For a system to have a value v of P (where P could be energy, position, etc.) with a non-minimal degree of determinacy D is to have stably differentiated quantum property D*-P to a non-minimal degree D* where D=D*. A system with a quantum property fully stably differentiated will have a determinate value of P.
On the other hand, indeterminacy and undifferentiation are related,
Note that according to this relation, we have multiple quantum properties concerning P, represented by quantum states and observables, that correspond to a non-maximally determinate value of P.139139139Note also that this relation doesn’t imply that undifferentiated quantum properties are more fundamental than indeterminate value properties.
@@
[81] enters into further details about this ontology. It may initially seem pedantic compared with the simpler ontology of flashes and observables. However, it captures more structure represented by quantum states (and decoherence) than the flashes. Systems don’t only have determinate values under interactions (which would be analogous to the flashes), they have quantum properties with different degrees of differentiation that change over time.
@@
In this appendix, I put into practice some of the above features of EnDQT to see how it can account for interference phenomena via a simple example.140140140This model is based on [92]. See [27] for a more extensive discussion of this framework.
@@
represents the number of photons (the particle number) in the channel $\mathrm{Ch}$, where each channel is associated with a radiation mode.
@@
After BS2, the system will interact with the detector D2, giving rise to a system having a 1 particle number determinate value during the interaction. Note that it is assumed that D2 is connected to an SDC.
Let’s clarify how interactions leading to a determinate value work by examining what happens when detector D3 is placed at B1 (see Figure 9). This detector interacts with the quantum system, annihilating the above interference phenomenon. I am going to adopt the same SDC subscripts convention that I have adopted in the last sections. The interactions at time t’ involving D3 (and omitting the interactions with D1 and D2) lead to the following state,
@@
where these interactions can be represented via decoherence models.141141141For simplicity, I will not analyze in detail decoherence in the Fock basis and assume that the Schrödinger picture is applicable. See [98, 66, 69] for a detailed account. I will also assume that a notion of spatiotemporal localization of particles arises during these interactions. See [38] for a survey of different options that consider particles as non-fundamental, but emergent. How does EnDQT interpret the above phenomenon? First, note that contrary to $\left|E_{1}(t)\right\rangle,\left|E_{0}(t)\right\rangle$ concerns the inexistence of the measurement signal. It will also mean that the measurement device interacted with a subsystem, giving rise locally to a 0 particle number determinate value (i.e., the vacuum).
@@
Let’s now consider instead the situation where the detectors are isolated from interacting with elements of an SDC, not belonging to an SDC as well. The quantum state $|Final(t^{\prime})\rangle$ isn’t anymore applicable to correctly represent the situation inside the lab. We would rather have
@@
Appendix E: From the Classical Markov Condition to the Classical Reichenbach Common Cause Principle
Let’s start with some key definitions.142142142This content is mainly based on [105, 101]. A directed graph $G$ over a collection of nodes $V$ is defined as the pair $\langle V,E\rangle$, where $E$ is a collection of directed edges. A path between nodes $X$ and $Y$ consists of a sequence of nodes that begin at $X$ and terminate at $Y$, in such a way that for each pair of consecutive nodes in the sequence, there exists a directed edge connecting them (the orientation of the edge is irrelevant). A node $X$ on a path is identified as a collider if, jointly with its adjacent nodes $Y$ and $Z$ on the path, it forms an inverted fork: $Y\rightarrow X\leftarrow Z$. Furthermore, nodes will be associated with random variables. Moreover, I will consider that there exists a directed path between nodes $X$ and $Y$, which I will denote as $X\rightsquigarrow Y$, if there is an arrow from $X$ to $Y$ ($X\rightarrow Y$) or if there is some node $Z$ such that $X\rightsquigarrow Z$ and $Z\rightarrow Y$.
Par(X) will be the set of parents of a node $X$, and it will include the nodes $Z$ in such a way that $Z\rightarrow X$. Childr(X), the set of $X$’s “children”, involves the nodes $Z$ such that $X\rightarrow Z$. Furthermore, I will establish the sets Anc(X) (referred to as “ancestors”) and Desc(X) (referred to as “descendants”) by substituting $\rightarrow$ with $\rightsquigarrow$ in the earlier definitions. It is important to note that a node is always considered its own ancestor and descendant, but it is never its own child or parent.
[…]
Figure 10: Example of a DAG
This definition can be shown to be equivalent to the one in the main text if it concerns a DAG. To see how it works, consider the DAG in the Figure 10. The CMC determines the following conditional independencies:
$B\perp\!\!\!\perp C,E\mid A$ (45)
$C\perp\!\!\!\perp B\mid A$ (46)
$D\perp\!\!\!\perp A,E\mid B,C$ (47)
$E\perp\!\!\!\perp A,B,D\mid C$ (48)
Then, we have the following definition of the principle of common cause:
The Classical Reichenbach Common Cause Principle (CRCCP) holds for a directed acyclic graph $G$ over $V$ and a probability distribution $P(V)$ if, whenever $A\rightsquigarrow B$ or $B\rightsquigarrow A$ or there is a $U\subseteq V$ such that $C\in U$ implies $C\rightsquigarrow A$ and $C\rightsquigarrow B$, and $A\perp\!\!\!\perp B\mid U$.
It is typically considered that the shortest proof of the connection between the CMC and the CMC appeals to $d$-separation. d-separation is not an intuitive concept. I will use the one in [48]. I have explained above the notion of a path. Now, consider $X$, $Y$, and $Z$ as disjoint subsets of $V$. $Z$ $d$-separates $X$ and $Y$ if and only if every path $\langle X_{1},\ldots,X_{k}\rangle$ from a variable in $X$ to a variable in $Y$ includes at least one variable $X_{i}$ such that either: (i) $X_{i}$ is a collider, and no descendant of $X_{i}$ (which includes $X_{i}$ itself) is in $Z$; or (ii) $X_{i}$ is not a collider, and $X_{i}$ is included in $Z$. A path that fulfills these criteria is considered to be blocked by $Z$. On the other hand, if $Z$ does not $d$-separate $X$ and $Y$, then $X$ and $Y$ are $d$-connected by $Z$. If $X$ and $Y$ are not d-separated by $Z$, then $X$ and $Y$ are $d$-connected by $Z$.
The $d$-separation is useful for our purposes due to the following feature ([101], pp. 17):
Given a directed acyclic graph $G$ over $V$ and $R,S,T\subseteq V$, $T$ $d$-separates $R$ and $S$ if and only if $R\perp\!\!\!\perp S\mid T$ for all probability functions that satisfy the Markov Condition with respect to $G$.
For instance, if we have knowledge that a certain DAG with a probability distribution fulfills the CMC, and we notice that different variables $X$ and $Y$ are $d$-separated by $\emptyset$, we are able to make the inference that $X$ and $Y$ are not correlated.
Let’s now turn to the proof that the CMC leads to the CRCCP. Let’s assume that the CMC holds for a DAG $G$ over $V$ and a $P(V)$. Let $A,B\in V$. Let’s assume that it is not the case that $A\rightsquigarrow B$ or $B\rightsquigarrow A$ or there is a $C\in V$ such that $C\rightsquigarrow A$ and $C\rightsquigarrow B$. Then, variables $A$ and $B$ are $d$-separated by $\emptyset$, because a collider has to be included in any path between them. Thus, in this case $A\perp\!\!\!\perp B$.
[…]